### v2

# A Data Generation Perspective to the Mechanism of In-Context Learning

Abstract
In-Context Learning (ICL) empowers Large Language Models (LLMs) with the capacity to learn in context, achieving downstream generalization without gradient updates but with a few in-context examples. Despite the encouraging empirical success, the underlying mechanism of ICL remains unclear, and existing research offers various viewpoints of understanding. These studies propose intuition-driven and ad-hoc technical solutions for interpreting ICL, illustrating an ambiguous road map. In this paper, we leverage a data generation perspective to reinterpret recent efforts and demonstrate the potential broader usage of popular technical solutions, approaching a systematic angle. For a conceptual definition, we rigorously adopt the terms of skill learning and skill recognition. The difference between them is skill learning can learn new data generation functions from in-context data. We also provide a comprehensive study on the merits and weaknesses of different solutions, and highlight the uniformity among them given the perspective of data generation, establishing a technical foundation for future research to incorporate the strengths of different lines of research.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) have revolutionized Natural Language Processing (NLP) (Achiam et al., 2023) and other relevant areas such as multimodality over vision and language (Liu et al., 2023a), accelerating numerous challenging research directions, e.g., AI agent (Durante et al., 2024), reasoning (Wei et al., 2022b), and story telling (Xie et al., 2023). These amazing applications are driven by LLMs’ emerging capacity of In-Context Learning (ICL), which can be formally defined as the new abilities that are not present in small models but arise in large models (Zhao et al., 2023). For instance, moral self-correction ability only occurs in LLMs of over 22B parameters (Ganguli et al., 2023).
Figure 1: An illustration of ICL example on the sentiment analysis. The upper instances with the background color of gray are the labeled in-context demonstrations, while the last line is the query for which LLMs infer the sentiment label.
In-Context Learning (ICL), a fundamental emerging capacity for many complicated abilities, is to utilize a few selected labeled demonstrations with the format (input, label) [^1] In this paper, we only consider classification tasks., before the query input, for making predictions in a few-/one-shot manner. An ICL example is illustrated in Figure 1.
Despite the empirical success of various ICL prompting strategies for downstream applications (Mavromatis et al., 2023; Ye et al., 2022), the fundamental mechanism of ICL remains unclear, leading to unexplainable ICL phenomenons, e.g., sensitivity to the sample order (Lu et al., 2021), robust to wrong input-label mapping (Min et al., 2022). Increasing attention has been paid to understanding the ICL from various perspectives. Nonetheless, this area is still in the infant stage. Due to the complexity of LLMs, most existing works only take one individual factor into account, e.g., pre-training data distribution (Chan et al., 2022), model scale (Wei et al., 2023), and in-context task difficulties (Raventos et al., 2023). Moreover, existing works involving the same factor may adopt different experimental settings (Yoo et al., 2022; Min et al., 2022), leading to potentially conflicted conclusions. Typically, Pan (2023) details ICL as two abilities: task recognition and task learning. The task recognition ability recognizes the underlying task of demonstrations and utilizes prior knowledge embedded in weights to solve the recognized task. In contrast, the task learning ability can learn new tasks, according to ICL demonstrations, unseen in the pre-training stage. Whereas, current understanding frameworks are intuitive and ad-hoc, hindering the progress of this area. A principled and systematic view of the ICL mechanism is highly desired.
In this paper, we propose the data generation perspective as a principled angle to comprehend existing research progress on the ICL mechanism. Such data generation perspective can be naturally found in both the pre-training stage and the ICL stage as the auto-regressive language model is generally utilized as the LLM backbone. Specifically, the pre-training objective is the next token prediction, aiming at modeling the data generation function of pre-training corpora; the objective of ICL inference is considered as a label generation process given the query inputs. Therefore, adopting the data generation perspective enables a unified framework through which we can cohesively address both the pre-training and the ICL stages, offering a holistic approach to understanding and leveraging the foundational mechanisms of LLMs.
Guided by the data generation perspective, we introduce a more principled and rigorous understanding framework on skill learning and skill recognition [^2] We refer to origin task learning as skill learning and origin task recognition as skill recognition, with an emphasis on the intrinsic capability of LLMs., distinguished by whether LLMs can learn a new data generation function in context. The skill learning ability learns a new data generation function in context. The new function is unseen in the pre-training stage and is (Bayesian) optimal to the given in-context demonstrations. The function learning statistical framework (Garg et al., 2022) is proposed particularly for the skill learning ability. The skill recognition ability selects one data generation function from the learned function class during pre-training. The Bayesian inference statistical framework (Xie et al., 2021) is generally utilized to understand the skill recognition ability. Notably, there exists a large gap between the two statistical frameworks, where the Bayesian inference framework utilizes the Hidden Markov Model (HMM) (Baum & Petrie, 1966) as the data generation function which is more complicated but more aligned with the real-world scenario. In contrast, the function learning framework commonly utilizes linear functions which are much simpler but distant to the real-world scenario.
Such a rigorous view also helps to ground the conceptual debate on whether the LLM can achieve intelligence into whether the LLM can learn new data generation functions in context, clearly unveiling the strengths and limitations of current LLM approaches to intelligent systems.
Our other contribution is, given the data generation perspective, we shed light on the transferability of two separated research lines of Bayesian inference and function learning. Specifically, we extend the Bayesian inference framework as a technical solution for skill learning, therefore opening the door to incorporate real-world cases, e.g., topics of corpora.
Organization: We introduce skill recognition with the Bayesian inference framework and skill learning with the function learning framework in Section 2 and 3, respectively. We then showcase the uniformity of the two frameworks via extending the original Bayesian inference framework for skill learning in Section 4. More empirical analysis on skill learning and skill recognition abilities are illustrated in Section 5. Future directions are discussed in Section 6.
Distinguish skill learning from skill recognition. The skill can be regarded as a data generation function, referring to the underlying hypothesis on the textual data generation. To determine whether the utilized skill is from the pre-training function class or is a new function, an empirical method is to validate whether LLMs can fit a set of data generated with a ground-truth function which is outside the pre-training function class.
Comparison with existing relevant literature. As far as we know, this paper is the first to provide a comprehensive discussion on existing studies about the mechanism of ICL and advocating a principled data generation perspective. This paper is different from existing surveys such as Dong et al. (2022); Zhao et al. (2023); Wei et al. (2022a).
Terminology. The prompt sequence of In-Context Learning consists of two parts: (1) The demonstrations illustrate as (input, label) pair, denoted as $(x_{i},y_{i})$; The demonstrations provide the basic description of the intended task. (2) The query is the test input after a few demonstrations. The ICL aims to provide the correct prediction for the query based on the in-context demonstrations and the knowledge of pre-trained LLM. The data generation function in this paper refers to the underlying hypothesis on textual data generation. It serves as the data assumptions in the theoretical understanding and the simulation data generator for the synthetic experimental analysis. Each data generation function obtained by the LLM can be recognized as a skill.
2 Skill Recognition
The skill recognition ability is to select the most proper data generation function from the function class obtained during pre-training, according to in-context demonstrations. A Bayesian inference framework (Xie et al., 2021) is introduced to explain skill recognition. The ICL inference can then be instantiated as a Bayesian inference process as follows:
$\small\begin{split}&p(\text{output}|\text{prompt})=\\ &\int_{\text{concept}}p(\text{output}|\text{concept},\text{prompt})p(\text{concept}|\text{prompt})d(\text{concept})\end{split}$ (1)
where $p(\text{output}|\text{prompt})$ is the conditional probability of the output generation given the prompt. It can be marginalized with the pre-training concepts and each concept corresponds to a pre-training data generation function. $p(\text{concept}|\text{prompt})$ is to locate the latent concept aligned with the in-context demonstrations, which will be the key focus in the ICL inference stage. After locating the aligned concept, $p(\text{output}|\text{concept},\text{prompt})$ is to utilize the selected data generate function for the output generation.
Such modeling method of latent concept is commonly adopted in the NLP domain as the language data is inherently compositional with unobserved latent concepts that are not explicitly present in the raw text, e.g., sentiments, topics, and syntactic structures (Chung et al., 2015; Zhou et al., 2020). Latent variable models can specify prior knowledge and structural dependency for language data which enjoys the characteristics of high compositionality. Deep latent variable models are popularly utilized to improve various tasks such as the alignment in statistical machine translation, topic modeling, and text generation (Kim et al., 2018; Fang et al., 2019; Wang et al., 2023).
However, there are various definitions of the aforementioned concept, all latent information that can help ICL to make the inference could be considered as a good choice of concept. We summarize the existing concept definitions as follows: (1) Xie et al. (2021) defines the concept as the transition matrix $\mathbf{\theta}$ of a Hidden Markov Model (HMM) (Baum & Petrie, 1966). The concept helps to state a transition distribution over observed tokens. A concrete example of the concept is the transition between name (Albert Einstein) $\to$ nationality (German) $\to$ occupation (physicist) ) in wiki bios. (2) Wang et al. (2023) simplifies the transition between tokens, modeled by HMM, with LDA topic models where each topic corresponds to one latent concept (Blei et al., 2003). However, the LDA topic model cannot capture the dependency between tokens. (3) Despite the above mathematical interpretations, Todd et al. (2023); Liu et al. (2023b) empirically establish the connection between the latent concept and the downstream task, e.g., supervised classification and question-answering, where the particular latent representation in the LLM can capture essential information about the task. The following discussions focus on the first concept definition (Xie et al., 2021) as the transition matrix of HMM.
Xie et al. (2021) utilizes the Bayesian inference framework to understand how the obtained pre-training data functions are activated by the in-context demonstrations. The key questions of this framework are: (1) In the pre-training stage, how the model obtained the latent concepts from the pre-training corpus; and (2) In the ICL inference stage, how the in-context demonstrations can locate the most relevant concept to generate the desired output for the query.
During the pre-training stage, the LLM can obtain various concepts from the large pre-training corpora if each pre-training document is generated from an individual HMM model. In such case, the next token prediction objective can converge if and only if the LLM can successfully generate the correct next token matching the HMM generation process dominated by the underlying concept. Notably, different documents can be generated from various concepts sampled from the concept set $\Theta$.
The ICL inference stage conducts an implicit Bayesian inference to locate the aligned pre-training concept $\theta^{*}\in\Theta$ with the optimal likelihood to generate the observed in-context demonstrations. The format of the prompt is shown below:
$\small\left[S_{n},x_{\text{test}}\right]=\left[x_{1},y_{1},o^{\text{del}},\ldots,x_{n},y_{n},o^{\text{del}},x_{\text{test}}\right]\sim p_{\text{prompt}}$ (2)
where $p_{\text{prompt}}$ is a data generation process implemented with HMM parameterized by the $\theta^{*}$. $x_{i}$, $y_{i}$ and $o^{\text{del}}$ are the input, label, and delimiter, respectively. The difficulty in locating $\theta^{*}$ is that it is of low probability for all the concepts to generate the in-context demonstrations. The key reason is that token transition patterns of the in-context demonstrations are in three types (1) the input to the label $x_{i}\to y_{i}$ (2) the label to the delimiter (3) the delimiter to the input. The latter two patterns hardly appear in the pre-training concepts due to different delimiter usage.
To address the above issue of low probability, the authors propose some strong assumptions that may not be feasible in real-world scenarios. We showcase one assumption as follows. It is more likely for the selected concept $\theta^{*}$ to transit to the delimiter hidden states while other concepts are unlikely to transit to the delimiter hidden states. Equipped with those assumptions, we are able to locate the aligned pre-training concept to implement Bayesian inference. Even though we cannot locate the aligned concept, Xie et al. (2021) provides the theoretical guarantee on the effectiveness of the ICL in such cases, where the ICL performance improves along with the in-context example length.
Inspired by the above Bayesian inference framework, more frameworks on understanding skill recognition are proposed, e.g., the PAC-Bayesian framework and Hopfield Network. Zhang et al. (2023b) further takes the architecture of transformers into consideration and analogies ICL inference to a Bayesian model averaging algorithm. Wies et al. (2023) propose a PAC-based generalization framework exhibiting satisfying generalization bound on the ICL where a transformer trained on multi-task can match the ICL performance of a transformer trained solely on the downstream task. Zhao (2023) analogizes the latent concept location as memory retrieval with the Hopfield Network. Nonetheless, all the existing studies are either theoretical or empirical based on synthetic data without considering the real-world scenario. It could be a promising future direction to investigate how the LLM is able to retrieve the concept with natural language explanations.
3 Skill Learning
The skill learning ability considers that the LLM can learn a new data generation function according to the in-context examples on the fly. For the interpretability of skill learning, we take the research line of function learning  [^3] We refer to algorithm learning as function learning with an emphasis on the approximated functions by algorithms and, in this way, it is easier to analyze ICL. for discussion. Specifically, pre-training is considered as a process to learn a class of functions that can fit the pre-training corpora, and the ICL inference is to learn a new data generation function to best fit ICL demonstrations.
The discussions on skill learning are organized as follows. In Section 3.1, we first provide a clear description of the function learning framework and illustrate its benefits and drawbacks. In Section 3.2, we investigate (1) the function fitting ability of ICL by examining whether ICL can learn new functions from a certain function class; and (2) the generalization ability of ICL. In Section 3.3, we explain the internal mechanisms why ICL can implement different learning algorithms.
3.1 The Function Learning Framework
The function learning framework is a prototype setting defining a new learning task. Denoting $\mathbf{x}\sim\mathcal{P}_{\mathcal{X}},\mathbf{x}\in\mathbb{R}^{d}$ where $\mathcal{P}_{\mathcal{X}}$ is a distribution, a function class $\mathcal{F}$ where for each $f\in\mathcal{F},f:\mathbb{R}^{d}\to\mathbb{R}$. Given a sequence ($\mathbf{x}_{1},\cdots\mathbf{x}_{i}$) ($i>1$) sampled from $\mathcal{P}_{\mathcal{X}}$ sequentially, and a sampled function $f\sim\mathcal{F}$, the learning objective aims to correctly predict $f(x_{i})$ based on the sequence $(\mathbf{x}_{1},f(\mathbf{x}_{1}),\cdots,\mathbf{x}_{i-1},f(\mathbf{x}_{i-1}),\mathbf{x}_{i})$ with both in-context examples and the query input $\mathbf{x}_{i}$.
[…]
where $\mathcal{L}$ is the loss function. $T_{\mathbf{\omega}}$ denotes the transformer model, $\mathbf{\omega}$ is the parameter of the transformer.
Notably, the model is pre-trained on an ICL objective rather than the original next-token prediction one. The function learning framework enables us to: (1) arbitrarily generate data with desired properties from the pre-defined function class $\mathcal{F}$; (2) clearly examine the function-fitting ability and the generalization of the skill learning in the ICL; and (3) utilize well-developed statistical learning theory to understand the ICL.
However, existing research on function learning is not sophisticated as it can not tell the characteristics of the learned function classes. For instance, some studies claim that ICL inference learns a gradient descent algorithm, but Fu et al. (2023a) shows a high-order version of the gradient descent algorithm. Besides, the function classes are learned implicitly during pre-training and ICL inference. Previous works prove this simulation by approximating the target algorithm, e.g., ridge regression, with a small training error gap between the ridge regression and self-attention networks (Bai et al., 2023; Li et al., 2023b). The simulation accuracy is not well discussed or compared. Moreover, this line of research has to consider both pre-training and ICL inference, and, due to the computational resource limitation, most works utilize transformers with less than 6 layers. Therefore, conclusions may not be generalizable to larger model scales.
3.2 Function-fitting and Generalization of ICL
In this subsection, we investigate the function fitting and the generalization behavior of the ICL. function fitting indicates to what extent transformers can learn unseen data generation functions in context based on the pre-training generation functions within the same class. Generalization, on the other hand, measures how well transformers can learn an unseen function best in the in-context demonstrations, and the in-context function is from one of the pre-training function classes, e.g. linear regression and decision tree.
To explore the function fitting, previous works validate if ICL inference can learn a similar function as the pre-training data generation function. Raventos et al. (2023) leverage separated data generation functions, sampled from a linear function class, for pre-training data and in-context demonstrations. With a smaller class of linear functions, transformers do not show the skill learning ability but the skill recognition one, and are a Bayesian optimal estimator, for the data generation function of in-context demonstrations, given the pre-training functions as a prior. Provided with a larger class of linear functions, all unseen functions in the linear function class can be implemented by skill learning, achieving a performance matching the optimal least squares estimator. Wu et al. (2023) further provides a theoretical explanation to support the above empirical observations.
Beyond the linear function class, Garg et al. (2022) observes that the ICL is expressive enough to implement more complicated functions when pre-training functions are also from the same function class, including sparse linear functions, two-layer neural networks, and decision trees. Bai et al. (2023); Fu et al. (2023b) propose theoretical explanations with a generalization bound between the prediction error of the toy transformer model and that of the target function, e.g. decision tree. However, there are still two essential questions unsolved: (1) Why the transformers suddenly obtain the skill learning ability with significant performance increase once the number of pre-training data generation functions reaches a certain threshold? (2) Why the learned data generation function of ICL demonstrations is from the same class as the pre-training data generation function?
The generalization of ICL is validated by comparing the ground-truth data generation function of in-context demonstrations and the predicted one. Assuming the pre-training data generation functions cover decision tree and linear function, the ground-truth data generation function of ICL demonstrations is a linear function. The ICL generalization is strong if and only if the predicted function of ICL demonstrations is a linear function.
Bai et al. (2023); Ahuja et al. (2023) indicate that the transformer can implement the Bayesian optimal selection which selects the best-fitting function, with the minimum description length, from all the function classes that appeared in the pre-training stage. Such Bayesian optimal selection helps a transformer pre-trained with multiple function classes reach comparable ICL performance as one pre-trained with the ground-truth function class. Notably, such Bayesian optimal on the synthetic dataset may not extend to the real-world More surprisingly, Yadlowsky et al. (2023) observes even if each pre-training instance is not generated from one individual function class. Instead, each instance is generated from a weighted sum of multiple functions from different function classes, e.g., $0.7f_{1}(x)+0.3f_{2}(x)$ where $f_{1}$ and $f_{2}$ are from different function classes. The same conclusion holds. Nonetheless, skill learning fails if the ground-truth function is out of the pre-training function class, then the transformer only demonstrates skill recognition ability (Yadlowsky et al., 2023).
Summary: Skill learning ability emerges if the function class of potential pre-training data generation functions is sufficiently large. For in-context demonstrations, ICL can learn a function that lies in the same function class of the pre-training data. The ICL would implement a Bayesian optimal selection to select the best-fitting function, aligned with ICL demonstrations generate function, from pre-training function classes.
3.3 The Internal Mechanisms of ICL
In this subsection, we further explore the mechanism of ICL by answering why it can learn an unseen function in context. Notably, there are two common assumptions in existing works: (1) The data generation functions for both pre-training data and in-context demonstrations are linear. (2) The toy transformer model is linearized by removing feed-forward layers and the softmax activation function in the attention layer. Ahn et al. (2023b) proposes indirect evidence that the linearized simplification does not hurt the precision of the vanilla transformer architecture, illustrating that the training dynamic of the linearized version is similar to that of the vanilla version.
Previous works analogize ICL to meta-learning (Finn et al., 2017). The pre-training stage corresponds to the outer-loop optimization, and the ICL inference stage is an instance of the inner-loop optimization, implementing fast adaptation on new novel tasks. Rather than a real inner gradient update, ICL inference mimics gradient update via a forward process with in-context demonstrations (Hubinger et al., 2019; von Oswald et al., 2023). Based on the dual view that a linear neural layer updated with gradient is equivalent to a linear attention forward process (Irie et al., 2022)Dai et al. (2022) proves the mathematical equivalence, illustrating the implicit gradient descent implementation with a linear attention. However, such an analogy is only limited to mathematical equivalence, it is not a straightforward answer to why ICL can learn a function since such an analogy overlooks all practical details on the gradient descent, including the choice of the learning objective, the pre-training weight, and the training data distribution.
Following-up research focuses on how the ICL forward process can implement different functions, given suitable assumptions on the transformer’s weights. Von Oswald et al. (2023) first demonstrate that the ICL on the single-layer transformer can implement one-step gradient descent with a linear regression objective. Bai et al. (2023) further show the ICL inference can implement ridge regression, least square, lasso, and even gradient descent on a two-layer Neural Network. Nonetheless, those strong assumptions about the attention weights may be not practically reasonable. For instance, Von Oswald et al. (2023) construct the key, query, value matrices $W_{K},W_{Q},W_{V}$ with $W_{K}=W_{Q}=\left(\begin{array}[]{cc}I_{x}&0\\ 0&0\end{array}\right),W_{V}=\left(\begin{array}[]{cc}0&0\\ W_{0}&-I_{y}\end{array}\right)$, where $I_{x}$ and $I_{y}$ are two different identity matrices and $W_{0}$ is the initialized parameters of the transformer model. Nonetheless, it is unclear why a pre-trained transformer would have such type of weights, and it has been reported not easily achieved in practice (Shen et al., 2023).
To extend the construction of attention weights aforementioned into a more practical setting, Zhang et al. (2023a); Mahankali et al. (2023); Ahn et al. (2023a) showcase the converged weights after the pre-training stage can simulate some learning algorithms during the ICL inference. However, the converged weights may not be precise enough since they are determined based on the loss landscape or gradient flow analysis. Typically, the ICL on the one-layer linear transformer can implement gradient descent (Von Oswald et al., 2023) or preconditioned gradient descent algorithm (Ahn et al., 2023a) given a linear regression objective. Given a two-layer transformer, ICL can implement a gradient descent with adaptive step size and special sparsity regularization (Ahn et al., 2023a). Moreover, Ahn et al. (2023a); Von Oswald et al. (2023) reveal that multiple-layered transformers can implement a GD++ algorithm. For larger-scale transformers, Akyürek et al. (2022) empirically illustrates that, instead of performing GD, large-scale transformers show emergent ability directly approximating the closed-form solution of ridge-regression, while there is still a gap on why this ability emerges as the model-scale increases.
Beyond the linear activation for attention heads, more recent research takes the softmax activation function into consideration. Von Oswald et al. (2023) demonstrate there exists a transformer that performs GD to solve more complicated nonlinear regression tasks. Li et al. (2023a); Ren & Liu (2023) identify the nonlinear regression task as the softmax regression and contrastive learning objective, respectively. Cheng et al. (2023) further take non-linear data generation functions into consideration, elucidating a transformer can implement gradient descent and converge to the Bayes optimal predictor. Guo et al. (2023) further studies a more challenging but practical setting of representation learning, in which predictions depend on inputs through the MLP. The theoretical evidence indicates that the ICL inference can implement ridge regression in context with the input of neural representations.
Summary: Existing research shows the mathematical equivalence between gradient descent and toy transformer models, working as proof that transformers implicitly implement gradient descent in context. However, this is supportive evidence for the function learning framework while there are no insights about how to select in-context demonstrations.
4 Insights on the Bayesian Inference and the Function Learning Framework
The skill recognition and skill learning abilities are comprehensively discussed with the example statistical frameworks of the Bayesian inference and function learning in Section 2 and 3, respectively. Instead of explaining one ability with one specific statistical framework, we exhibit the potential that both frameworks can be easily utilized to understand the mechanism of both abilities. Such extension enables the future mechanism analysis to select the suitable analysis framework, by referring to their strengths and weaknesses. The function learning framework provides an elegant description of the data generation process with more comprehensive conclusions. However, it is over-simplified with an unclear relevance to the real-world scenario. The Bayesian inference framework provides a more concrete and detailed description of the data generation process through an HMM model, e.g., the delimiter is taken into consideration, while the theoretical analysis on the role of delimiters is hard since it requires several assumptions over statistical modeling.
The function learning framework can be easily extended to understand skill recognition by simply replacing the data generation function from a mixture of HMMs with a mixture of linear functions. In this section, we focus on how to utilize the Bayesian inference framework to model the mechanism of skill learning. We first show that the original function learning framework for the skill learning ability also implements an implicit Bayesian optimal selection in Section 4.1. We then extend the Bayesian inference framework to learn new in-context data generate functions in Section 4.2, therefore the Bayesian inference framework can also be a technical solution for skill learning.
4.1 Bayesian Selection in the Function Learning Framework
The Bayesian perspective can be found in the function learning framework originally utilized for the skill learning mechanism. Typically, we illustrate the underlying Bayesian selection in the function learning framework, indicating the intrinsic connection between the two statistical frameworks. According to Ahuja et al. (2023), the transformers pre-trained on the data generated from diverse function classes exhibit improved function-fitting ability across all the pre-training function classes. To identify the best-fit solution among the whole function class, the function selection process implements a Bayesian optimal selection. More details can be found in Section 3.2. Notably, instead of the original Bayesian inference framework only selecting the pre-training data generation functions, the function selection scope is enlarged, including all the unseen functions from the same function class with the pre-training functions.
4.2 Extending the Bayesian Inference Framework for Skill Learning
We then illustrate the possibility of extending the Bayesian inference framework to understand the skill learning mechanism to capture new data generation functions from the in-context data via relaxing the particular assumption. One important assumption in the Bayesian inference framework (Xie et al., 2021) is that all ICL demonstrations should be generated with the same latent concept. Nonetheless, this strong assumption may not be held in practice. For instance, one demonstration sample discusses the topic of sociology but another one is relevant to cardiology, the data generation function for these two domains should be rather different. Inspired by the high compositionality nature of language data, Hahn & Goyal (2023) came up with an information-theoretic bound showing that ICL performance can be improved given more unique compositional structures in pre-training data, therefore skill learning ability can appear by combining compositionality structures, in pre-training data, to infer the data generation function of ICL demonstrations.
Empirical evidence shows that, given an input-label pair of two semantically unrelated concepts, e.g., mapping sports to animals, Rong (2021); Wei et al. (2023) still observe a satisfactory performance with the increasing model scale, indicating that the LLM can retrieve multiple concepts and combine them as a new data generation process. Feng & Steinhardt (2023) interpret the combination with a binding mechanism with an internal function vector to recognize the input feature and bind it to the corresponding label.
Swaminathan et al. (2023) proposes another way to extend the existing Bayesian framework for skill learning via replacing the original HMM model into the clone-structured causal graph (CSCG) (George et al., 2021; Dedieu et al., 2019). The major difference is that the CSCG considers a learnable emission matrix, which determines the probability of observing a particular output given each hidden state in the model. A relevant transition matrix as the concept is retrieved, similar to the Bayesian inference (Xie et al., 2021). The hidden states for each token can then be obtained given the particular relevant template. The LLM then learns the suitable emission matrix, providing the best-fit mapping from the hidden states to the observed token.
5 Empirical Investigation On Skill Recognition and Skill Learning
In this section, we exhibit more empirical analyses revolving around skill recognition and skill learning abilities. In contrast to the mechanism analysis that focuses on whether the ICL can learn new in-context data generation functions or not, empirical evidence in this section indicates that it is highly likely that LLMs exhibit both skill recognition and skill learning abilities of various levels, instead of an all-or-nothing conclusion. We first discuss how the LLM jointly obtains both abilities during the pre-training stage in Section 5.1. Specifically, the origin of both abilities is determined by the pre-training data distribution (Chan et al., 2022) and the model scale (Wei et al., 2023; Pan, 2023). We then investigate how the LLM effectively utilizes the obtained abilities during the ICL inference stage in Section 5.2. Typically, the LLM exhibits varying degrees of usage on those two abilities according to tasks with different difficulties. The unique strengths and weaknesses of each ability are shown in Section 5.3.
5.1 Origin of Skills
In this subsection, we carefully examine how well the LLM obtains the skill learning and the skill recognition abilities during the pre-training stage, with a focus on the impact of the pre-training data distribution and model scale. Roughly speaking, the skill recognition ability is easy to achieve while the skill learning ability develops much slower and only emerges when the model scale is sufficiently large.
Analyses are first conducted focusing on how those abilities are developed along the pre-training procedure. (Bietti et al., 2023) observe that the skill recognition ability is obtained early in the pre-training procedure, while the skill learning ability is developed much later. However, Singh et al. (2023) shows that the obtained skill learning ability gradually vanishes after over-training and is replaced by the skill recognition ability. Such observation indicates that skill learning is a transient ability that may disappear when the model is over-trained rather than a persistent one which can be kept once obtained. The reason can be attributed to the pre-training data distribution (Chan et al., 2022) where the task learning ability degrades if the pre-training data follows a uniform, i.i.d distribution. Nonetheless, such degradation may not happen when the pre-training data follows a properly skewed Zipfian distribution. Chan et al. (2022) further emphasizes that the skill learning ability emerges when the pre-training data meets the following properties: (1) Skewed rank-frequency distributions: Dynamic contextual meaning does not uniform across data, instead, only a few meanings dominate with the long tail of other infrequent meanings. (2) Burstiness: Dynamic contextual meaning is not uniform across time, but appears in clusters. The reason why ICL ability can be obtained on such data distribution remains unclear. A potential explanation could be that the pre-training weight can only obtain the head meaning frequently appears while the long tail knowledge can only be obtained via ICL.
Analyses are then conducted with a focus on the impact of the model scale. Pan (2023) illustrates that the skill recognition ability can be found across LLMs with different scales. In contrast, LLMs obtain better skill learning ability along with an increasingly larger scale. Similar observations can be found in (Wei et al., 2023) that the LLM can learn the flipped input-label mapping and override pre-training knowledge when the model scale is sufficiently large. (Fu et al., 2023c) provides the potential explanation for the above phenomenon where the good skill recognition ability serves as a necessity for developing the skill learning ability.
5.2 Application of Skills
After the LLM obtained the skill learning and skill recognition abilities during pre-training, we then investigate how the model utilizes both abilities for achieving satisfactory downstream task performance during the ICL inference stage. Overall, the behavior of the LLM is more consistent with the skill recognition mechanism on difficult tasks while observations aligned with skill learning are more common to see on easy tasks.
Empirical analyses are conducted on the well-trained LLM, focusing on the ICL behavior on downstream tasks with various difficulties. Typically, we examine whether the model behavior aligns with the skill recognition ability or the skill learning one via the performance sensitivity on corrupting in-context data with incorrect input-label mapping. If the LLM takes advantage of the skill learning ability more, the LLM can learn the corrupted in-context mapping, leading to performance degradation compared with the origin setting. In contrast, if the LLM follows the skill recognition ability more, the LLM should be robust to the correctness of the input-label mapping, since the skill recognition ability only implements the pre-training data generation function with correct input-label mapping. Min et al. (2022) first observes that the corrupted mapping does not necessarily lead to the overall performance degradation, indicating an overall skill recognition behavior. Instead of examining the overall performance across tasks, Yoo et al. (2022) conducts a more careful evaluation of each task individually where the ICL shows different behaviors on tasks with different difficulties. The relatively easy tasks exhibit performance degradation on the wrong input-label mapping while the robust performance appears on those difficult tasks. Such observation indicates that the skill learning ability is more applicable to relatively easy tasks while the skill recognition ability dominates on the difficult ones.
5.3 Advantage and Disadvantage of Skills
Considering the intricate interplay of both abilities on different tasks, we further illustrate the strengths and weaknesses inherent in each ability. Skill learning ability can obtain new knowledge from the in-context data, and even over-ride the pre-training knowledge. It provides an easy way to update the knowledge on the specific application without requiring computationally heavy fine-tuning. Such ability has been successfully utilized in different LLM applications, e.g. model editing with ICL (Zheng et al., 2023). Nonetheless, the skill learning ability may fail as it can be easily distracted by irrelevant context (Shi et al., 2023). The failure reason found in (Tang et al., 2023) is that the input-label mapping favors the shortcut as the model scale increases. Skill recognition ability is insensitive to the new in-context pattern leading to the failure on the specification-heavy task (Peng et al., 2023) while it exhibits robustness to the incorrectness of label-demonstrations and other in-context noise (Webson & Pavlick, 2021). Based on the above discussion, we suggest a careful evaluation of LLM about each ability and select a desired one for the downstream task.
6 Future Directions
How the skill learning ability emerges during pre-training. The emergence of the skill learning ability can be partially attributed to the skewed rank-frequency distribution of pre-training corpora (Chan et al., 2022), and Reddy (2023) highlights the role of the induction head (Olsson et al., 2022), a particular attention head which explicitly searches for a prior occurrence of the current token in-context and copying the suffix as prediction. Moreover, the function class-based analysis (Raventos et al., 2023) illustrates that the transition from skill recognition to skill learning only happens given diverse enough tasks in pre-training corpora. It is interesting to explore how these factors collaboratively influence the emergence of skill learning.
Why ICL can only learn data generation function appeared during the pre-training? In section 3, we provide a comprehensive discussion on what function can be learned in context. Observations indicate that ICL can only learn the function within the pre-training data generation function class. Nonetheless, the causality of the pre-training data generation function to ICL remains unclear. Garg et al. (2022) proposes the research question as Can we train a model to in-context learn a certain function class but overlooks the effect of the pre-training data generation function class. Once we have a certain clue about causality, we can leverage the skill-learning ability in a more controllable and safe manner.
Extending existing findings to other capabilities of LLMs. Despite the very original ICL being for classification tasks, more capabilities are observed, e.g., step-by-step reasoning ability (Wei et al., 2022b) for reasoning and self-correction (Ganguli et al., 2023). A critical question is how we can extend the understanding frameworks introduced in this paper, particularly the data generation perspective, to more complicated LLMs’ capabilities. Some pioneering research has been done; Prystawski & Goodman (2023) extends the Bayesian inference framework to understand the effectiveness of the CoT prompt. Kadavath et al. (2022) focuses on the self-evaluation prompt showing that LLMs can accurately examine the correctness of their statements. We believe the introduced data generation perspective and two main understanding frameworks on ICL serve us the milestone to explore more intrinsic capacities of LLMs.
7 Conclusion
In this study, we introduce a novel data generation perspective and use it to understand the underlying mechanism driving the current success of ICL. We primarily focus on understanding the LLM’s ability of skill learning and skill recognition, and investigating whether or not ICL inference is capable of learning new data generation functions in context. Given the increasing significance of LLMs, enhancing our understanding of these underlying processes is crucial. Our work makes a step forward in this direction, hopefully, it will stimulate further investigation of this important topic.